# Open-Vocabulary Camouflaged Object Segmentation

### v2

@@ 5 Experiments
| Model | VLM | Feature Backbone | Text Prompt | cS | cF | cMAE | cF | cE | cIoU |
@@
| SimSeg [49] | CLIP-ViT-B/16 [37] | ResNet-101 [20] | Learnable [60] | 0.128 | 0.105 | 0.838 | 0.112 | 0.143 | 0.094 |
| OVSeg [27] | CLIP-ViT-L/14 [37] | Swin-B [30] | [18] | 0.341 | 0.306 | 0.584 | 0.325 | 0.384 | 0.273 |
| ODISE [47] | CLIP-ViT-L/14 [37] | StableDiffusionv1.3 [40] | [17] | 0.409 | 0.339 | 0.500 | 0.341 | 0.421 | 0.302 |
| SAN [48] | CLIP-ViT-L/14 [37] | ViT Adapter | [18] | 0.414 | 0.343 | 0.489 | 0.357 | 0.456 | 0.319 |
| CAT-Seg [10] | CLIP-ViT-L/14 [37] | Swin-B [30] | [37] | 0.430 | 0.344 | 0.448 | 0.366 | 0.459 | 0.310 |
| FC-CLIP [52] | CLIP-ConvNeXt-L [9] | — | [18] | 0.374 | 0.306 | 0.539 | 0.320 | 0.409 | 0.285 |
@@
| SimSeg [49] | CLIP-ViT-B/16 [37] | ResNet-101 [20] | Learnable [60] | 0.098 | 0.071 | 0.852 | 0.081 | 0.128 | 0.066 |
| OVSeg [27] | CLIP-ViT-L/14 [37] | Swin-B [30] | [18] | 0.164 | 0.131 | 0.763 | 0.147 | 0.208 | 0.123 |
| ODISE [47] | CLIP-ViT-L/14 [37] | StableDiffusionv1.3 [40] | [17] | 0.182 | 0.125 | 0.691 | 0.219 | 0.309 | 0.189 |
| SAN [48] | CLIP-ViT-L/14 [37] | ViT Adapter | [18] | 0.321 | 0.216 | 0.550 | 0.236 | 0.331 | 0.204 |
| CAT-Seg [10] | CLIP-ViT-L/14 [37] | Swin-B [30] | [37] | 0.185 | 0.094 | 0.702 | 0.110 | 0.185 | 0.088 |
| FC-CLIP [52] | CLIP-ConvNeXt-L [9] | — | [18] | 0.124 | 0.074 | 0.798 | 0.088 | 0.162 | 0.072 |
@@
| SimSeg [49] | CLIP-ViT-B/16 [37] | ResNet-101 [20] | Learnable [60] | 0.053 | 0.049 | 0.921 | 0.056 | 0.098 | 0.047 |
| OVSeg [27] | CLIP-ViT-L/14 [37] | Swin-B [30] | [18] | 0.024 | 0.046 | 0.954 | 0.056 | 0.130 | 0.046 |
| ODISE [47] | CLIP-ViT-L/14 [37] | StableDiffusionv1.3 [40] | [17] | 0.187 | 0.119 | 0.700 | 0.211 | 0.298 | 0.167 |
| SAN [48] | CLIP-ViT-L/14 [37] | ViT Adapter | [18] | 0.275 | 0.202 | 0.612 | 0.220 | 0.318 | 0.189 |
| CAT-Seg [10] | CLIP-ViT-L/14 [37] | Swin-B [30] | [37] | 0.181 | 0.106 | 0.719 | 0.123 | 0.196 | 0.094 |
| FC-CLIP [52] | CLIP-ConvNeXt-L [9] | — | [18] | 0.080 | 0.076 | 0.872 | 0.090 | 0.191 | 0.072 |
| Ours | CLIP-ConvNeXt-L [9] | — | CamoPrompts | 0.579 | 0.490 | 0.337 | 0.520 | 0.615 | 0.443 |
| Model | cS | cF | cMAE | cF | cE | cIoU | |
@@
5.1 Implementation Details
Dataset Settings. As mentioned in Sec. 3, the proposed dataset is divided according to classes into two disjoint subsets, i.e., and . The former contains 14 classes for training and the latter contains the remaining 61 classes for testing. The proposed dataset itself is provided with only images and masks. We use the typical monocular depth estimation method DPT [38] to obtain the depth map for training, while the edge map is generated by dilating and eroding operations. To avoid information leakage during testing, we use depth and edge maps only in the training phase. And these generated depth maps and edge maps will be made publicly available with the dataset.
Model Settings. Following previous settings in [52, 49, 48], the pre-trained CLIP is frozen during training, and the remaining parameters are learnable and are randomly initialized. The AdamW [31] optimizer with the learning rate of 3e-6, weight decay of 5e-4, batch size of 4, and training epoch of 30 is used to optimize model parameters. Some basic data augmentations including random flipping, rotating, and color jittering, are introduced to preprocess training data. The input data is bilinearly interpolated to 384 384 during training and inference.
Evaluation Protocol. To reasonably evaluate the performance of OVCOS, we modify the metrics in the original CSU task [13, 8] to cS, cF, cMAE, cF, cE, and cIoU which follows the common settings in the OVSIS field [49, 27, 48, 10, 52, 61] and takes into account both classification and segmentation.
@@ 5.1 Implementation Details
| 0 |
"<class>" w/o MAP based on ground truth |
0.538 |
@@
| 4 |
templates from [37] |
0.686 |
| 5 |
templates from [18] |
0.682 |
@@
| 6.9 |
6.7, 6.8 |
0.703 |
| 6.10 |
6.1 - 6.9 i.e., CamoPrompts |
0.704 |
5.2 System-level Comparison
5.3 Analysis and Ablation Study
Importance of Modules. To specifically analyze the effect of different components, we evaluate their performance in Tab. 3. As can be seen, the proposed modules all show positive gains. Both the semantic guidance in the class-aware decoding and the structure enhancement from auxiliary tasks of depth and edge estimation consistently boost the final performance. The ablation comparison also demonstrates that explicit guidance of the spatial and contour information is important for the detection of camouflaged objects. Besides, in Tab. 3, we also give the ideal results of our CLIP-driven framework where is treated as . Although our method exhibits leading performance compared to existing methods as shown in Tab. 2, there is still a long way to go to solve this problem. At the same time, the current ideal performance is still far from the limit, suggesting that future breakthroughs in this field may require more powerful paradigms.
Importance of Iterative Refinement. The top-down iterative refinement strategy significantly improves the OVCOS performance as shown in Tab. 3. When the number of iterations is 2, our algorithm obtains the best OVCOS performance. As increases, there is no further improvement in performance, so it is set to 2 by default. In addition, the correlation guidance from the output space plays an important role, and introducing the object-aware representation also has positive gains.
Importance of CamoPrompts. Our prompt template set CamoPrompts, which takes task attributes into account, shows better performance in Fig. 8. To further understand the influence of different templates on the semantic embedding, we calculate the Hausdorff distance between training and testing class labels in the embedding space, as shown in Fig. 8. The figure presents an interesting phenomenon that those templates with better classification performance tend to reduce the distance, which inspires further explorations for more effective prompt engineering.
@@ 5.3 Analysis and Ablation Study
Model Complexity. We compare the proposed method with the CLIP-based competitors [49, 27, 47, 48, 10, 52] in terms of the number of trainable and total parameters, and FLOPs. To be fair, we test all methods following the setting of their original inference settings. As can be seen from Fig. 9, our method has fewer trainable parameters () and less computational complexity (), which is superior to these competitors.

### v3

@@ 5 Experiments
5.1 Implementation Details
Dataset Settings. As mentioned in Sec. 3, the proposed dataset is divided according to classes into two disjoint subsets, i.e., and . The former contains 14 classes for training and the latter contains the remaining 61 classes for testing. The proposed dataset itself is provided with only images and masks. We use the typical monocular depth estimation methods [51, 26] to obtain the depth map for training, while the edge map is generated by dilating and eroding operations. We introduce depth and edge maps only in the training phase and they are not required during inference, which also avoids information leakage during testing. So, our depth and edge maps are labor-free and training-only, and directly generated by existing algorithms. Such a design can help our model get rid of the dependence on them in real-world applications. Our structure enhancement based on depth and edge information belongs to a deeper exploration of this task. These generated depth maps will be made publicly available with the dataset.
Training Details. Following previous settings in [70, 64, 63], the pre-trained CLIP is frozen during training, and the remaining parameters are learnable and are randomly initialized. The AdamW [39] optimizer with the learning rate of 3e-6, weight decay of 5e-4, batch size of 4, and training epoch of 30 is used to optimize model parameters. Some basic data augmentations including random flipping, rotating, and color jittering, are introduced to preprocess training data. The input data and the output logits are bilinearly interpolated to during training and inference.
Evaluation Protocol. To reasonably evaluate the performance of OVCOS, we modify the metrics in the original CSU task [13, 8] to cSm, cF, cMAE, cFβ, cEm, and cIoU which follows the common settings in the OVSIS field [64, 33, 63, 10, 70, 89] and takes into account both classification and segmentation.
5.2 System-level Comparison
| Model | VLM | Feature Backbone | Text Prompt | cSm | cF | cMAE | cFβ | cEm | cIoU |
@@ 5.2 System-level Comparison
| SimSeg21 [64] | CLIP-ViT-B/16 [50] | ResNet-101 [20] | Learnable [88] | 0.128 | 0.105 | 0.838 | 0.112 | 0.143 | 0.094 |
| OVSeg22 [33] | CLIP-ViT-L/14 [50] | Swin-B [37] | [18] | 0.341 | 0.306 | 0.584 | 0.325 | 0.384 | 0.273 |
| ODISE23 [62] | CLIP-ViT-L/14 [50] | StableDiffusionv1.3 [53] | [17] | 0.409 | 0.339 | 0.500 | 0.341 | 0.421 | 0.302 |
| SAN23 [63] | CLIP-ViT-L/14 [50] | ViT Adapter | [18] | 0.414 | 0.343 | 0.489 | 0.357 | 0.456 | 0.319 |
| CAT-Seg23 [10] | CLIP-ViT-L/14 [50] | Swin-B [37] | [50] | 0.430 | 0.344 | 0.448 | 0.366 | 0.459 | 0.310 |
| FC-CLIP23 [70] | CLIP-ConvNeXt-L [9] | — | [18] | 0.374 | 0.306 | 0.539 | 0.320 | 0.409 | 0.285 |
@@
| SimSeg21 [64] | CLIP-ViT-B/16 [50] | ResNet-101 [20] | Learnable [88] | 0.098 | 0.071 | 0.852 | 0.081 | 0.128 | 0.066 |
| OVSeg22 [33] | CLIP-ViT-L/14 [50] | Swin-B [37] | [18] | 0.164 | 0.131 | 0.763 | 0.147 | 0.208 | 0.123 |
| ODISE23 [62] | CLIP-ViT-L/14 [50] | StableDiffusionv1.3 [53] | [17] | 0.182 | 0.125 | 0.691 | 0.219 | 0.309 | 0.189 |
| SAN23 [63] | CLIP-ViT-L/14 [50] | ViT Adapter | [18] | 0.321 | 0.216 | 0.550 | 0.236 | 0.331 | 0.204 |
| CAT-Seg23 [10] | CLIP-ViT-L/14 [50] | Swin-B [37] | [50] | 0.185 | 0.094 | 0.702 | 0.110 | 0.185 | 0.088 |
| FC-CLIP23 [70] | CLIP-ConvNeXt-L [9] | — | [18] | 0.124 | 0.074 | 0.798 | 0.088 | 0.162 | 0.072 |
@@
| SimSeg21 [64] | CLIP-ViT-B/16 [50] | ResNet-101 [20] | Learnable [88] | 0.053 | 0.049 | 0.921 | 0.056 | 0.098 | 0.047 |
| OVSeg22 [33] | CLIP-ViT-L/14 [50] | Swin-B [37] | [18] | 0.024 | 0.046 | 0.954 | 0.056 | 0.130 | 0.046 |
| ODISE23 [62] | CLIP-ViT-L/14 [50] | StableDiffusionv1.3 [53] | [17] | 0.187 | 0.119 | 0.700 | 0.211 | 0.298 | 0.167 |
| SAN23 [63] | CLIP-ViT-L/14 [50] | ViT Adapter | [18] | 0.275 | 0.202 | 0.612 | 0.220 | 0.318 | 0.189 |
| CAT-Seg23 [10] | CLIP-ViT-L/14 [50] | Swin-B [37] | [50] | 0.181 | 0.106 | 0.719 | 0.123 | 0.196 | 0.094 |
| FC-CLIP23 [70] | CLIP-ConvNeXt-L [9] | — | [18] | 0.080 | 0.076 | 0.872 | 0.090 | 0.191 | 0.072 |
| Ours | CLIP-ConvNeXt-L [9] | — | CamoPrompts | 0.579 | 0.490 | 0.336 | 0.520 | 0.616 | 0.443 |
| Model | cSm | cF | cMAE | cFβ | cEm | cIoU | |
@@
5.3 Analysis and Ablation Study
Better Baseline Model. Interestingly, our baseline model in Tab. 3 outperforms existing methods in Tab. 2. We attribute this to the fact that the overly complex feature pipeline of the conventional OVSIS architecture is not well adapted to this task. And our framework, from hyperparameter setting to structural evolution, is explored for the proposed task and data. Thus, we actually provide a more effective and robust baseline setup for OVCOS.
Importance of Modules. To specifically analyze the effect of different components, we evaluate their performance in Tab. 3. As can be seen, the proposed modules all show positive gains. Both the semantic guidance in the class-aware decoding and the structure enhancement from auxiliary tasks of depth and edge estimation consistently boost the final performance. The ablation comparison also demonstrates that explicit guidance of the spatial and contour information is important for the detection of camouflaged objects and their positive gains are subject to the effective fusion design. We replace the proposed structure enhancement (SE) component with the vanilla addition fusion with a similar number of parameters, which results in detrimental effects. Besides, in Tab. 3, we also give the ideal results of our CLIP-driven framework where is treated as . Although our method exhibits leading performance compared to existing methods as shown in Tab. 2, there is still a long way to go to solve this problem. At the same time, the current ideal performance is still far from the limit, suggesting that future breakthroughs in this field may require more powerful paradigms as discussed in Appendix 0.C.
Importance of Iterative Refinement. The top-down iterative refinement strategy significantly improves the OVCOS performance as shown in Tab. 3. When the number of iterations is 2, our algorithm obtains the best OVCOS performance. As increases, there is no further improvement in performance, so it is set to 2 by default. In addition, the object-aware representation from the output space plays an important role, and introducing the correlation guidance also has positive gains.
Importance of CamoPrompts. Our prompt template set CamoPrompts, which takes task attributes into account, shows better performance in Sec. 5.3. To further understand the influence of different templates on the semantic embedding, we calculate the Hausdorff distance between training and testing class labels in the embedding space, as shown in Fig. 9. The figure presents an interesting phenomenon that those templates with better classification performance tend to reduce the distance, which inspires further explorations for more effective prompt engineering.
@@ 5.3 Analysis and Ablation Study
| 0 | "<class>" w/o MAP based on ground truth | 0.538 |
@@
| 4 | templates from [50] | 0.686 |
| 5 | templates from [18] | 0.682 |
@@
| 6.9 | 6.7, 6.8 | 0.703 |
| 6.10 | 6.1 - 6.9 i.e., CamoPrompts | 0.704 |
@@
Model Complexity. We compare the proposed method with the CLIP-based competitors [64, 33, 62, 63, 10, 70] in terms of the number of trainable and total parameters, and FLOPs. To be fair, we test all methods following the setting of their original inference settings. As can be seen from Sec. 5.3, our method has fewer trainable parameters () and less computational complexity (), which is superior to these competitors.
@@ 6 Conclusion
Acknowledgements. We sincerely thank these authors of the public datasets [14, 84, 65, 2, 8] used in this work: Deng-Ping Fan, Ge-Peng Ji, Guolei Sun, Ming-Ming Cheng, Jianbing Shen, and Ling Shao for COD10K [14], Yunfei Zheng, Xiongwei Zhang, Feng Wang, Tieyong Cao, Meng Sun, and Xiaobing Wang for CPD1K [84], Jinyu Yang for PlantCamo [65], Pia Bideau and Erik Learned-Miller for CAD [2], and Xuelian Cheng, Huan Xiong, Deng-Ping Fan, Yiran Zhong, Mehrtash Harandi, Tom Drummond, and Zongyuan Ge for MoCA-Mask [8]. It is their efforts that underpin this new exploration of camouflaged object segmentation.
Appendix 0.A Model Details
0.A.1 Details of and
In the proposed model, the visual encoder and embedding layer together are used to extract the high-level embedding corresponding to the object of interest in the input image. The two independent sub-networks are split from the visual network of CLIP [50, 9]. contains all the feature encoding layers for extracting the multi-scale image features. And corresponds to the final high-dimensional projection layer, which is used to convert the high-level image feature into the visual embedding vector .
0.A.2 Multi-scale Image Features
The multi-scale image features is from the four stages of ConvNeXt [38] with different output resolutions, respectively. While the feature is obtained by up-sampling the feature by bilinear interpolation.
0.A.3 Class Embedding
In the proposed algorithm, for each image input, the textual encoder needs to extract text embedding for all class texts. However, due to the nature of the algorithm design, these text embeddings are shared in each iteration, so they can be pre-computed in the inference to save inference cost.
0.A.4 Details of Classification
During the inference, the class prediction is generated from the pair-wise correlation matrix after the softmax operation, where is from the multiplication between the normalized visual embedding and textual embedding, i.e. and . Besides, in our experiments, introducing classification supervision on top of the existing form interferes with the training process of the model, which brings about significant performance degradation, with relative reductions of 8.6% and 15.7% for cSm and cF, respectively. Therefore, we do not consider classification loss during training, and is only used to construct the top-down iterative guidance.
Appendix 0.B Dataset Details
0.B.1 Class Semantic Similarity
To analyze the semantic similarity between the relabelled classes, we compute the path similarity between the classes based on the Open English WordNet [41] as shown in Fig. 11. Specifically, when is the length of the shortest path between two classes, the path similarity score is:
The score ranges between 0.0 and 1.0, where the higher the score is, the more similar the two classes are. The score is 1.0 when a class is compared to itself, and 0.0 when there is no path between the two classes (i.e., the path distance is infinite). As shown in Fig. 11(b), the semantic similarity between classes in our class set is very low. Most of them lie around 0.1, while the maximum is only 0.5. Such low similarity can better alleviate the complexity due to class semantic similarity during open vocabulary evaluation.
0.B.2 Class Hierarchy Relationships
In Fig. 12, we use the alluvial graph to show the class relationships at different levels, including super, base, and sub-classes. The sub-classes shown here include class names with clearer meanings preserved from the original data and class names after initial manual correction. The base classes represent the class names obtained after careful manual filtering and merging, which was used in all the experiments in this paper. The super classes generalize the base classes from a broader perspective.
Appendix 0.C Limitations and Future Works
0.C.1 Class Embedding Setting
In our proposed algorithm, the class embedding setting follows the existing OVSIS methods. Although our iteration process does not impose much computational burden due to our caching mechanism, there is still a computational complexity associated with the number of classes. This is not ideal for practical applications in open vocabulary scenarios. More flexible and efficient class embedding designs are still worth exploring.
0.C.2 CLIP-based Architecture
The transfer application of the CLIP in downstream dense prediction tasks is limited by its pre-training form and the camouflage scenes that we focus on may be more affected. Some work [33, 63, 10] attempts to further finetune CLIP, but the finetuning strategy still needs to be designed more carefully due to the potential disruption to the CLIP’s open vocabulary ability. And it also suggests that there is room for further improvement. Besides, as mentioned in the main text, the current ideal performance of the CLIP-based architecture is still far from the limit, which suggests that future breakthroughs in this field may require more powerful paradigms.
0.C.3 Data Scale
Although the number of finely labeled samples in our proposed dataset is over ten thousand, the data scale is still smaller than existing large datasets such as COCO-Stuff [4] and ADE20K [85]. So there is still a need to collect more data, especially for classes with fewer samples. We can resort to rich web images, which also may lead to significant manual labeling costs. In addition, the data synthesis technique has been demonstrated in some recent work [66, 67] to greatly facilitate the performance of semantic image segmentation tasks. This technique based on object masks and textual descriptions in existing datasets, may bring some new insights. But this also needs to address the ensuing interference problem.
0.C.4 Class Scale
Open-vocabulary segmentation as a hot topic, currently focuses on how to use the open-vocabulary capability of VLMs to segment objects with unseen classes. More data classes will indeed facilitate the development of OVCOS. However, the imperceptibility of camouflaged objects poses a great challenge for further expansion. This is the focus of our future work.
0.C.5 Prompt
The importance of prompt engineering for visual language modeling can be reflected in the existing literature [50, 18, 64, 33, 70, 62, 10, 63] and the experiments in this paper. The prompt forms used in this paper rely heavily on manual design, which is still limited by the knowledge of the prompter. More automated prompt-generation strategies may be needed in the future, which deserve more attention and exploration.
Appendix 0.D Dataset Copyright
We have investigated the copyright information of these data sources, and they are currently now widely used in the CSU field [15], and available for non-commercial academic use. Much of existing work [34, 58] provides only new annotations and an index of the original data rather than the data itself. And users can download original data from the sources. Following these existing community practices, for the proposed OVCamo dataset, we list the data links provided by the original authors in the documentation. We also provide the class annotations we created and a detailed description of the way the data is organized in OVCamo. In addition, we thank the contributors of the relevant datasets in the our acknowledgements.